whats the cone surface area if the r is 5 the slanted hiegt is 9 and 3.14 is used for pi
Answer by lordshubham(4)

There are Two types of Surface Areas for Cone here are both:
TOTAL SURFACE AREA OF CONE
= π*r(r+l)
= 3.14*5(5+9)
= 15.7(14)
= 219.8cm^2
Curved Surface Area of Cone= πrl
= 3.14*5*9
=141.3 cm^2